# JoyVASA: Portrait and Animal Image Animation with Diffusion-Based Audio-Driven Facial Dynamics and Head Motion Generation

### v4

@@ 3.3.1 Audio Conditioned Diffusion Model
In the diffusion model, we aim to sample predicted motions $\hat{X}^{0}$ from noisy observations $X^{t}$, conditioned on the corresponding audio features $A$, where $t\in T$ denotes the discrete time steps in the diffusion process. To improve temporal consistency in motion generation, we adopt a strategy that includes both the previous and current windows of features in the model’s training. Specifically, at each diffusion step $t$, the network takes as input both the past and present speech features $[A_{-W_{pre}}:A_{W_{cur}}]$, the previous motion parameters $X_{-W_{pre}}^{0}$, and the current noisy motion parameters $X_{W_{cur}}^{t}$, which are sampled from $q(X^{t}_{0:W_{cur}}|X^{0}_{0:W_{cur}})$. The denoising network then generates the clean motion samples, as described by the following equation:
@@ 3.4 Inference
The inference pipeline is illustrated in Figure 1. Given a reference image, we first extract the corresponding 3D facial appearance feature $f_{face}$ using the appearance encoder $E_{app}$, and the learned motion information $[x_{c},R_{s},\delta_{s},t_{s}]$ using the motion encoder $E_{mnt}$. For the input speech, the audio features are initially extracted using the wav2vec2 encoder. The audio-driven motion sequences $\hat{X}^{0}$ are then sampled using the diffusion model trained in the second stage in a sliding window fashion. Using the canonical source keypoints $x_{c}$ and the sampled target motion sequences $\hat{X}^{0}$, the target keypoints $x_{d}$ are computed as shown in equation 1. Finally, the 3D facial appearance feature is warped based on the source and target keypoints and rendered by a generator to produce the final output video.
@@ 5.2 Comparison with existing methods
Method IQA(%) $\uparrow$ VQA(%) $\uparrow$ Sync-C $\uparrow$ Sync-D $\downarrow$ FVD-25$\downarrow$ Smooth(%) $\uparrow$

### v5

@@ 3.3.1 Audio Conditioned Diffusion Model
In the diffusion model, we aim to sample predicted motions $\hat{X}^{0}$ from noisy observations $X^{t}$, conditioned on the corresponding audio features $A$, where $t\in T$ denotes the discrete time steps in the diffusion process. To improve temporal consistency in motion generation, we adopt a strategy that includes both the previous and current windows of features in the model’s training. Specifically, at each diffusion step $t$, the network takes as input both the past and present speech features $A_{-W_{pre}}:A_{W_{cur}}]$, the previous motion parameters $X_{-W_{pre}}^{0}$, and the current noisy motion parameters $X_{W_{cur}}^{t}$, which are sampled from $q(X^{t}_{0:W_{cur}}|X^{0}_{0:W_{cur}})$. The denoising network then generates the clean motion samples, as described by the following equation:
@@ 3.4 Inference
The inference pipeline is illustrated in Figure 1. Given a reference image, we first extract the corresponding 3D facial appearance feature $f_{face}$ using the appearance encoder $E_{app}$, and the learned motion information $x_{c},R_{s},\delta_{s},t_{s}]$ using the motion encoder $E_{mnt}$. For the input speech, the audio features are initially extracted using the wav2vec2 encoder. The audio-driven motion sequences $\hat{X}^{0}$ are then sampled using the diffusion model trained in the second stage in a sliding window fashion. Using the canonical source keypoints $x_{c}$ and the sampled target motion sequences $\hat{X}^{0}$, the target keypoints $x_{d}$ are computed as shown in equation 1. Finally, the 3D facial appearance feature is warped based on the source and target keypoints and rendered by a generator to produce the final output video.
@@ 5.2 Comparison with existing methods
[…]